# CLARITY: Contextual Linguistic Adaptation and Accent Retrieval for Dual-Bias Mitigation in Text-to-Speech Generation

### v1

@@ 3.2 Experimental Results and Bias Analysis
1. RAAP and Contextual Linguistic Text Adaptation: Fig. 3 shows the RAAP accuracy (left) and the adapted scores generated by GPT-5 and LLaMA models (right). RAAP achieves nearly 100% accuracy for most accents in terms of accent and gender retrieval, while age prediction is less accurate but acceptable, given its known difficulty. For the LLM-as-judge scores, we observe that the standard text already achieves reasonable accent scores, with all scores above 6. After adaptation using GPT-4o-mini, four accents (GB, MY, SG, and US) show improved adaptation scores. Interestingly, these accents correspond to locations where the language of daily communciation tends to be English. A possible explanation is that for these speakers, the model had higher confidence in modifying the text, while the training corpus distribution likely reinforces this behaviour, which GPT-4 effectively captures. In contrast, LLaMA shows decreased accent scores after adaptation, although we doubt these results since the judge is GPT-5, not LLaMA, and continue to evaluate both adapted texts.
@@
2. Comparision of Baselines and CLARITY: Fig. 4 plots accent accuracy and NISQA scores for generated speech across various systems, along with the averaged scores. CosyVoice2 and ParlerTTS achieve very low accent accuracy (8.75% and 6.14%), likely due to training sets lacking accented speech. Per-accent results reveal higher scores for accents better represented in training data. CLARITY with GPT-adapted text achieves the best overall accuracy, though ES, JP, and PT remain below 50%, showing the difficulty of these accents. In terms of NISQA, ParlerTTS averages 4.67, while CLARITY_GPT achieves 4.36, outperforming CosyVoice2 (3.16) and even ground truth on some accents.
@@
3. Ablation Study of CLARITY: Table 1 presents the ablation study, showing that accent-score–guided prompt selection is highly effective, while adding text adaptation further improves accent accuracy. All variants achieve NISQA above 4.21, rising to 4.31 with RAAP and text adaptation, confirming the robustness of CLARITY. Among CLARITY variants, using GPT-adapted text performs best, while with LLaMA-adapted text is competitive, though its adapted text scores are lower (Fig. 3), indicating potential mismatches from using the GPT-5 as judge. Comparison of LLaMA and GPT for text selection does not improve results, suggesting that GPT-5 scores may not be the optimum [^11] Using multiple LLM-as-Judge models, e.g., GPT-5 and LLaMA-4, may reduce such mismatches, we leave this for future work.. Since GPT-adapted CLARITY performs best in the objective test, we adopt this setting for CLARITY in the following experiments.
Fig. 6: FDR comparison.
Fig. 7: Predicted accent distributions for baseline vs. proposed systems. TP = True Positive, FP = False Positive; correctly predicted accents are shadowed.
@@
Subjective Evaluation: Fig. 5 shows violin plots of the listening test results. For naturalness, ParlerTTS performs best on IN, while CLARITY outperforms both alternatives on CN and GB. SG results are similar across systems, with CosyVoice2 slightly worse. This aligns with the NISQA scores in Fig. 4. CLARITY demonstrates superior accent consistency with reduced bias across all accents. Both CLARITY and ParlerTTS maintain gender consistency, whereas CosyVoice2 mimics prompt speech rather than following user instructions. Overall, CLARITY best aligns with user instructions.
@@
Bias Analysis: The FDR across all accent groups in Fig. 6 shows that CLARITY systems with higher FDRs achieve greater fairness than the baselines. The distribution of true and false predictions across accents for CosyVoice2 and CLARITY in Fig. 7 indicates that CosyVoice2 (left) heavily biases toward US and CA accents (large dark-green and dark-red bars), often predicting them regardless of the real/true accent. CLARITY reduces this bias (right), decreasing US/CA dominance and increasing correctly predicted segments (shadowed), especially for GB and IN, resulting in more accurate and balanced accent distributions. This is particularly evident for GB and IN, where correctly predicted segments are much larger in CLARITY than in the baseline. Moreover, CLARITY achieves better subjective results in Fig. 5 for accent consistency and overall match scores, further confirming the effectiveness of the proposed method in mitigating accent bias.
4 Conclusion
@@
The proposed CLARITY framework demonstrates that jointly mitigating accent and linguistic bias in instruction-guided TTS leads to measurable gains in inclusivity and authenticity. Across twelve English accents, CLARITY achieves the highest accent accuracy, competitive naturalness, and significantly improved fairness compared to state-of-the-art baselines. Human listening tests confirm reduced bias perception and stronger alignment with target accents. Although validated on twelve accents, the framework is inherently backbone-agnostic and can be easily extended to a wider range of accents, with future work focusing on multilingual adaptation and deeper accent–linguistic alignment.
@@ Appendix A Appendix
A.1 Experimental Protocol
A.1.1 Accent Pool Data Generation
Table 2: The statistics of Accent Pool (AESRC)
@@ A.1.1 Accent Pool Data Generation
Table 3: The statistics of Accent Pool (SEAME)
@@
Table 4: Overview of scenarios and metadata coverage in the curated instruction dataset.
@@
Fig. 8: Distributions of gender (left) and age (right) across all generated implicit instructions.
We evaluate our proposed method on twelve English accents. The accent pool is composed of accents from two datasets. Tables 2 and 3 summarize the statistics of the accent pool, including the number of speakers, utterances, gender distribution, and age range. Specifically, ten accents are drawn from the AESRC dataset, which was introduced in the Interspeech 2020 Accented English Speech Recognition Challenge [shi2021accented]. This dataset comprises approximately 180 hours of speech spanning the ten English accents. Each speaker reads sentences covering common conversational topics as well as human–computer interaction commands. For our experiments, we selected a subset of 52,614 utterances, balanced by gender and covering a broad age range. For each speaker, up to 100 longest utterances were retained to ensure consistent recording quality and sufficient duration per sample. The selected subset includes the following accents: Canadian (CA), Chinese (CN), Spanish (ES), British (GB), Indian (IN), Japanese (JP), Korean (KR), Portuguese (PT), Russian (RU), and American (US).
The remaining two accents, Malaysian (MY) and Singaporean (SG), are represented by 22,108 utterances selected from the SEAME dataset [lyu2010seame], a large-scale, code-switching English–Mandarin corpus collected from spontaneous conversations in Malaysia and Singapore. It contains speech from both Malaysian and Singaporean speakers, with detailed accent and speaker annotations. To construct our subset, we extracted English-dominant utterances from both conversational and interview recordings, retaining only those labeled as English or code-switched (EN and CS) in the transcript metadata. Purely Mandarin (ZH) segments were excluded. Each utterance was required to contain at least 5 words, with a maximum of 100 utterances per speaker, prioritizing longer segments. After extraction, only speakers with English accent confidence scores greater than 0.9 were retained. These filtering criteria yielded a balanced selection of Malaysian and Singaporean speakers across both genders and an age range of 18 to 33 years.
A.1.2 Instruction Curation
@@
Table 4 summarizes the three scenarios and their corresponding metadata diversity. All implicit instructions were generated automatically using the GPT-4 language model [openai2023gpt4]. For each pair of a base sentence and sampled metadata, the model was prompted to provide paralinguistic guidance that specifies how the sentence should be spoken. The generation prompt was structured as follows:
@@
Fig. 9: Pipeline from user instructions to LLM-as-a-judge
A.1.3 Adapted Text Generation
@@
Format your output as a JSON object like this: { ”accent”: ”Speaker accent. Choose from the above pool. If there is no exact match, choose the closest available option.”, ”language”: ”language in the phrase to be spoken, e.g. EN”, ”age”: ”exact_age — [range_start, range_end] – use exact age if given, e.g. 25; if approximate age only, infer a 10-year range, e.g. [20, 30].”, ”gender”: ”M — F”, ”tone”: ”e.g. soft, angry, romantic, etc.”, ”emotion”: ”optional emotion, e.g. love, sadness, happiness”, ”additional_context”: ”any inferred intent or style, e.g. persuasive, affectionate, instructional” }
@@
Return a JSON object with only the localized ”text” field: {{ ”text”: ”<<localized text here>>.” }}
@@
IMPORTANT: Reward localisation only when it feels natural and contextually appropriate. The presence of a local particle or slang word does not automatically make a sentence better. If the expression feels forced, awkward, or out of place, the score should be reduced accordingly. Format your output as a JSON object like this: {{ ”score”: <<a list of scores in order of the samples>>, ”reason”: ”<<a list of short reasonings in order of the samples>>” }}
Fig. 10: Combined visualization of per-accent accuracy (left) and confusion matrix (right) on the validation set.
Fig. 11: Average NISQA for each accent category. Higher scores indicate better perceptual quality.
A.2 Additional Results on Accent Pool data
A.2.1 Accent Recognition Accuracy
To evaluate the performance of the accent preservation on the generated speech, we fine-tuned ECAPA-TDNN model, pretrained on the CommonAccent corpus [zuluaga2023commonaccent], for 12-class accent classification (CA, CN, GB, IN, JP, KR, MY, PT, RU, SG, ES, US) using the AESRC2020 and SEAME datasets. The model extracts 80-dimensional log-Mel filterbank features (25 ms window, 10 ms hop) followed by mean-variance normalization. The ECAPA-TDNN backbone (embedding dimension = 192, $\sim$6.2 M parameters) was kept frozen, and a single linear layer with 12 output units was fine-tuned. We used a batch size of 8, AdamW optimizer with a learning rate of $1\times 10^{-4}$, and AAM loss as the objective. The model was trained for 15 epochs with a random seed of 42 on an NVIDIA RTX 4070 GPU. The dataset was randomly divided into 80% for training and 20% for validation.
@@ A.2.1 Accent Recognition Accuracy
A.2.2 Speech Quality Evaluation
@@
Table 5: Ablation study of RAAP, comparing text similarity using standard user-provided text or LLM-adapted text.
@@
Fig. 12: Subjective evaluation results for GPT-based methods using either adapted or standard text for text similarity in prompt speech selection.
Fig. 13: Female evaluators
Fig. 14: Male evaluators
Fig. 15: Predicted accent distributions for baseline vs. proposed systems. TP = True Positive, FP = False Positive; correctly predicted accents are shadowed.
A.3 Additional Results on Generated Speech
A.3.1 Results on Retrieved Prompt Speech
@@
The objective results are presented in Table 5, where adapted text yields slightly lower accent accuracy for both LLaMA and GPT, while achieving comparable NISQA scores. In the subjective listening test, GPT-based methods using either adapted or standard text were evaluated. The violin plots in Figure 12 show that for SG, the results are almost identical between standard and adapted text; for CN, a similar trend is observed except that MOS with standard text achieves slightly better performance; while for GB and IN, GPT with standard text as the basis for text similarity calculation tends to obtain higher scores across all perceptual evaluation metrics.
A.3.2 Subjective Analysis
@@
A.3.3 Bias Analysis
Statistical Analysis - Binomial Test: to specifically test for the baseline model’s bias toward US and CA accents, we use a Binomial test. The event in this context is: a true non-US/CA accent sample being incorrectly predicted as a US/CA accent by the model. Our Null hypothesis (H_0H\_0) is as follows: The probability of the model predicting a non-US/CA accent as a US/CA accent is equal to the random chance probability (with 12 total accents, the random chance is $1/12\approx 0.083$). This represents no model bias. On the other hand, if the probability of the model predicting a non-US/CA accent as a US/CA accent is significantly higher than the random chance probability, it indicates that the model has a bias.
Table 6: Binomial test: biases towards US and CA accents comparision for CosyVoice2 and CLARITY

### v2

@@ III-B Objective Evaluation
1. RAAP and Contextual Linguistic Text Adaptation: Fig. 3 shows the RAAP accuracy (left) and the adapted scores generated by GPT-5 and LLaMA models (right). RAAP achieves nearly 100% accuracy for most accents in terms of accent and gender retrieval, while age prediction is less accurate but acceptable, given the known challenges related to this [18]. For the LLM-as-judge scores, we observe that the standard text already achieves reasonable accent scores, with all scores above 6. After adaptation using GPT-4o-mini, four accents (GB, MY, SG, and US) show improved adaptation scores. Interestingly, these accents correspond to locations where the language of daily communication tends to be English. A possible explanation is that for these speakers, the model had higher confidence in modifying the text, while the training corpus distribution likely reinforces this behaviour, which GPT-4 effectively captures. By contrast, LLaMA shows decreased accent scores after adaptation, possibly because the judge is GPT-5, not LLaMA. We evaluate both adapted text outputs further.
@@
2. Comparision of Baselines and CLARITY: Fig. 4 plots accent accuracy and NISQA scores for generated speech across various systems, along with the averaged scores. CosyVoice2 and ParlerTTS achieve very low accent accuracy (8.75% and 6.14%), likely due to training sets lacking accented speech. Looking in detail at individual accent results reveal generally higher scores for accents better represented in training data. CLARITY with GPT-adapted text achieves the best overall accuracy, though ES, JP, and PT remain below 50%, indicating the difficulty of mitigating outputs for these accents. In terms of NISQA, ParlerTTS averages 4.67, while CLARITY_GPT achieves 4.31, easily outperforming CosyVoice2 (3.16). Further testing revealed that CosyVoice2 works much better when it is prompted with a known-correct target accent example, but that was was not possible in our zero-shot evaluation scenario.
@@
3. Ablation Study of CLARITY: Table I presents the ablation study, showing that accent-score–guided prompt selection (second column) is highly effective, while incorporating text adaptation (third column) further improves accent accuracy. All variants achieve a NISQA above 4.21, rising to 4.31 with RAAP and text adaptation, confirming the robustness of CLARITY. Among CLARITY variants, using GPT-adapted text performs best, while LLaMA-adapted text systems are competitive albeit scoring slightly lower (confirming Fig. 3). Results may indicate a potential mismatche from using the GPT-5 as judge [^10] Using multiple LLM-as-judge models, e.g., GPT-5 and LLaMA-4, may reduce such mismatches – we leave this for future work.. Since GPT-adapted CLARITY performs best in the objective test, we adopt this setting for the following experiments.
Figure 6: FDR comparison (higher is better), averaged over all accent types, for the three systems, indicating that CLARITY outperforms both CosyVoice2 and ParlerTTS in terms of ensuring accent fairness.
Figure 7: Predicted accent distributions for two systems. The predicted accents are plotted in colour for the target accents along the x-axis. Correctly predicted accents are shadowed for CosyVoice2 (left) and CLARITY (right).
III-C Subjective Evaluation:
@@
Listening test naturalness results are presented as violin plots in Fig. 5 for four example dialects of English (Singapore, India, Great Britain and China, clockwise from top left), comparing scores obtained from CosyVoice2, ParlerTTS and CLARITY. It is clear that CLARITY peforms best for CN and GB whereas ParlerTTS performs best on IN. The SG results are similar across systems. The results align well with the NISQA scores in Fig. 4, confirming that CLARITY demonstrates superior accent consistency with reduced bias across all accents. In fact both CLARITY and ParlerTTS maintain gender consistency, whereas CosyVoice2 mimics prompt speech rather than following user instructions. Overall, the results shown that CLARITY best aligns with user instructions.
III-D Bias Analysis:
@@
Fairness Discrepany Rate: We compute the Fairness Discrepancy Rate (FDR) [3, 6] for all accent groups, as plotted in Fig. 6, revealing that CLARITY achieves higher FDRs than the baselines. To explore further, we consider the distribution of true and false positive predictions of all accents for CosyVoice2 and CLARITY generated speech in Fig. 7 with the predicted accent type represented as colour for real accents along the x-axis, excluding US/CN for visibility. CosyVoice2 (left) clearly biases heavily toward generating US and CA accents (large dark-green and dark-red bars), often regardless of the requested and purported accent being generated. CLARITY reduces this bias (right) by decreasing US/CA dominance and increasing the correctly predicted segments (shadowed). This is especially noticeable for GB, IN and SG, where correctly predicted segments are much larger in CLARITY than in the baseline. This tends to explain why CLARITY achieved better subjective results in Fig. 5 for those languages, in terms of accent consistency as well as overall match scores.
TABLE II: Binomial test: biases towards US and CA accents comparision for CosyVoice2 and CLARITY
Binomial Text To specifically test for the baseline model’s bias toward US and CA accents, we use a Binomial test. The event in this context is: a true non-US/CA accent sample being incorrectly predicted as a US/CA accent by the model. Our Null hypothesis (H_0) is as follows: The probability of the model predicting a non-US/CA accent as a US/CA accent is equal to the random chance probability (with 12 total accents, the random chance is $1/12\approx 0.083$). This represents no model bias. On the other hand, if the probability of the model predicting a non-US/CA accent as a US/CA accent is significantly higher than the random chance probability, it indicates that the model has a bias. The P-value  [^11] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.binomtest.html of US and CA accents are shown in Table VII. The P-value indicates the result of the Binomial test, where a very small P-value (typically $p\leq 0.05$) means the observed proportion is significantly higher than the random chance, thus confirming the existence of a statistically significant bias. We can see that CosyVoice2 has a extreme and consistent bias towards US/CA accents across all tested non-US/CA accents. CLARITY_GPT demonstrates a much lower and more variable bias compared to CosyVoice2 which reduced overall bias.
IV Conclusion
@@
The proposed CLARITY framework demonstrates that jointly mitigating accent and linguistic bias in instruction-guided TTS leads to measurable gains in fairness and accent authenticity. The authors believe that this leads to better inclusivity of those models, as well as mitigating lack of target cultural and linguistic knowledge of users when generating for minority or less well known accents. When evaluated on twelve variant English accents, CLARITY was shown to achieve the highest accent accuracy, produce highly competitive naturalness, and significantly improve fairness, compared to state-of-the-art frameworks. Human listening tests confirmed these results in reporting reduced bias perception and stronger alignment to target accents for a sunset of four tested target accents. The presented CLARITY framework is inherently backbone-agnostic, meaning it can be used with any zero-shot instruction-guided TTS system, and it can be easily extended to a wider range of accents. Future work will focus more on this multilingual adaptation as well as more advanced mixture-of-LLM judging of accent-linguistic alignment.
@@ Appendix A Appendix
A-A Experimental Protocol
A-A1 Accent Pool Data Generation
We evaluate our proposed method on twelve English accents. The accent pool is composed of accents from two datasets. Tables III and IV summarize the statistics of the accent pool, including the number of speakers, utterances, gender distribution, and age range. Specifically, ten accents are drawn from the AESRC dataset, which was introduced in the Interspeech 2020 Accented English Speech Recognition Challenge [16]. This dataset comprises approximately 180 hours of speech spanning the ten English accents. Each speaker reads sentences covering common conversational topics as well as human–computer interaction commands. For our experiments, we selected a subset of 52,614 utterances, balanced by gender and covering a broad age range. For each speaker, up to 100 longest utterances were retained to ensure consistent recording quality and sufficient duration per sample. The selected subset includes the following accents: Canadian (CA), Chinese (CN), Spanish (ES), British (GB), Indian (IN), Japanese (JP), Korean (KR), Portuguese (PT), Russian (RU), and American (US).
The remaining two accents, Malaysian (MY) and Singaporean (SG), are represented by 22,108 utterances selected from the SEAME dataset [12], a large-scale, code-switching English–Mandarin corpus collected from spontaneous conversations in Malaysia and Singapore. It contains speech from both Malaysian and Singaporean speakers, with detailed accent and speaker annotations. To construct our subset, we extracted English-dominant utterances from both conversational and interview recordings, retaining only those labeled as English or code-switched (EN and CS) in the transcript metadata. Purely Mandarin (ZH) segments were excluded. Each utterance was required to contain at least 5 words, with a maximum of 100 utterances per speaker, prioritizing longer segments. After extraction, only speakers with English accent confidence scores greater than 0.9 were retained. These filtering criteria yielded a balanced selection of Malaysian and Singaporean speakers across both genders and an age range of 18 to 33 years.
TABLE III: The statistics of Accent Pool (AESRC)
@@ A-A1 Accent Pool Data Generation
TABLE IV: The statistics of Accent Pool (SEAME)
@@
TABLE V: Overview of scenarios and metadata coverage in the curated instruction dataset.
@@
Figure 8: Distributions of gender (left) and age (right) across all generated implicit instructions.
A-A2 Instruction Curation
@@
Table V summarizes the three scenarios and their corresponding metadata diversity. All implicit instructions were generated automatically using the GPT-4 language model. For each pair of a base sentence and sampled metadata, the model was prompted to provide paralinguistic guidance that specifies how the sentence should be spoken. The generation prompt was structured as follows:
@@
Figure 9: Pipeline from user instructions to LLM-as-a-judge
A-A3 Adapted Text Generation
@@
Format your output as a JSON object like this:
”accent”: ”<Speaker accent. Choose from the above pool. If there is no exact match, choose the closest available option.>”,
”language”: ”<language in the phrase to be spoken, e.g. EN>”,
”age”: ”exact_age — [range_start, range_end] – use exact age if given, e.g. 25; if approximate age only, infer a 10-year range, e.g. [20, 30].”,
”gender”: ”M — F”,
”tone”: ”<e.g. soft, angry, romantic, etc.>”,
”emotion”: ”<optional emotion, e.g. love, sadness, happiness>”,
”additional_context”: ”<any inferred intent or style, e.g. persuasive, affectionate, instructional>” }
@@
Return a JSON object with only the localized ”text” field:
{{ ”text”: ”<localized text here>.” }}
@@
IMPORTANT: Reward localisation only when it feels natural and contextually appropriate. The presence of a local particle or slang word does not automatically make a sentence better. If the expression feels forced, awkward, or out of place, the score should be reduced accordingly. Format your output as a JSON object like this:
”score”: <a list of scores in order of the samples>,
”reason”: ”<a list of short reasonings in order of the samples>
Figure 10: Combined visualization of per-accent accuracy (left) and confusion matrix (right) on the validation set.
Figure 11: Average NISQA for each accent category. Higher scores indicate better perceptual quality.
A-B Additional Results on Accent Pool data
A-B1 Accent Recognition Accuracy
To evaluate the performance of the accent preservation on the generated speech, we fine-tuned ECAPA-TDNN model, pretrained on the CommonAccent corpus, for 12-class accent classification (CA, CN, GB, IN, JP, KR, MY, PT, RU, SG, ES, US) using the AESRC2020 and SEAME datasets. The model extracts 80-dimensional log-Mel filterbank features (25 ms window, 10 ms hop) followed by mean-variance normalization. The ECAPA-TDNN backbone (embedding dimension = 192, $\sim$6.2 M parameters) was kept frozen, and a single linear layer with 12 output units was fine-tuned. We used a batch size of 8, AdamW optimizer with a learning rate of $1\times 10^{-4}$, and AAM loss as the objective. The model was trained for 15 epochs with a random seed of 42 on an NVIDIA RTX 4070 GPU. The dataset was randomly divided into 80% for training and 20% for validation.
@@ A-B1 Accent Recognition Accuracy
A-B2 Speech Quality Evaluation
@@
TABLE VI: Ablation study of RAAP, comparing text similarity using standard user-provided text or LLM-adapted text.
@@
Figure 12: Subjective evaluation results for GPT-based methods using either adapted or standard text for text similarity in prompt speech selection.
Figure 13: Female evaluators
Figure 14: Male evaluators
Figure 15: Predicted accent distributions for baseline vs. proposed systems. TP = True Positive, FP = False Positive; correctly predicted accents are shadowed.
A-C Additional Results on Generated Speech
A-C1 Results on Retrieved Prompt Speech
@@
The objective results are presented in Table VI, where adapted text yields slightly lower accent accuracy for both LLaMA and GPT, while achieving comparable NISQA scores. In the subjective listening test, GPT-based methods using either adapted or standard text were evaluated. The violin plots in Figure 12 show that for SG, the results are almost identical between standard and adapted text; for CN, a similar trend is observed except that MOS with standard text achieves slightly better performance; while for GB and IN, GPT with standard text as the basis for text similarity calculation tends to obtain higher scores across all perceptual evaluation metrics.
A-C2 Subjective Analysis
@@
A-C3 Bias Analysis
Statistical Analysis - Binomial Test
To specifically test for the baseline model’s bias toward US and CA accents, we use a Binomial test. The event in this context is: a true non-US/CA accent sample being incorrectly predicted as a US/CA accent by the model. Our Null hypothesis (H_0) is as follows: The probability of the model predicting a non-US/CA accent as a US/CA accent is equal to the random chance probability (with 12 total accents, the random chance is $1/12\approx 0.083$). This represents no model bias. On the other hand, if the probability of the model predicting a non-US/CA accent as a US/CA accent is significantly higher than the random chance probability, it indicates that the model has a bias.
TABLE VII: Binomial test: biases towards US and CA accents comparision for CosyVoice2 and CLARITY